OpenStudy (samara8954):

Glenn and Helen are playing ping pong. For every 3 points that Glenn scores, Helen scores 7 points. If Glenn scores 9 points, how many points does Helen score?

OpenStudy (samara8954):

189 21 10 63 END STUDY SESSION Copyright © 2014 Edmentum - All rights reserved.

OpenStudy (samara8954):

@lizzyrockefeller

OpenStudy (samara8954):

@dtuniyants

OpenStudy (mathmate):

Use fractions: \(\large \frac{Helen}{Glenn}=\frac{7}{3}=\frac{how~many}{9}\)

OpenStudy (samara8954):

what is that

OpenStudy (samara8954):

would the answer be 21

OpenStudy (mathmate):

That's correct. Can you show me how you got it?

OpenStudy (samara8954):

yes or no

OpenStudy (samara8954):

Daniel watches 1 hour of television for every 2 hours of homework he does each week. If he did 6 hours of homework last week, how many hours of television did he watch?

OpenStudy (samara8954):

12 6 4 3

OpenStudy (samara8954):

im guessing four or 3 i think 3

OpenStudy (anonymous):

yes ur correct its 3

OpenStudy (anonymous):

@Samara8954